# Gueter Josmy Faure1    Jia-Fong Yeh1    Min-Hung Chen2    Hung-Ting Su1 Winston H. Hsu1,3    Shang-Hong Lai4

### v2

@@ 4.3 Ablation Studies
Table 7: Performance comparison of MA-LMM with and without ECO / SeTR  integration
@@
We evaluate the impact of our Episodic COmpressor (ECO) as plug-ins to the existing MA-LMM model (He et al., 2024) by replacing MA-LMM’s memory bank. ECO is designed to efficiently process long video sequences while preserving temporal coherence and narrative structure. The results in Table 4.3 demonstrate substantial improvements with ECO integration with an accuracy increase by 3.4%. Moreover, ECO demonstrates superior efficiency compared to MA-LMM’s memory bank, almost halving the overall inference latency.
@@
To demonstrate the effectiveness and efficiency of our proposed ECO, we conduct a comparative analysis against two strong existing compression techniques: ToMe (Bolya et al., 2022) and MA-LMM (He et al., 2024) in fig. 5. We calculate the inference time for each model on the MovieChat-1k dataset. our model achieves the highest accuracy (78.6%) among all models, outperforming MA-LMM by 5.3% and ToMe by a substantial 13.8%. Despite its superior accuracy, HERMES maintains competitive inference speed, comparable to ToMe and notably faster than MA-LMM with a significant 46% reduction in inference time compared to the latter. These results highlight our model’s ability to deliver state-of-the-art accuracy while maintaining efficiency.
@@
Our experiments on the MovieChat-1k (zero-shot) and Breakfast (fully-supervised) datasets reveal compelling insights into the optimal configuration of ECO (Figure 4) and SeTR (Figure 4). For ECO, we discover that an episodic memory size of 20 consistently yields peak performance across both datasets, achieving a 78.6% accuracy on MovieChat-1k and a 95.2% on Breakfast. This sweet spot balances comprehensive video representation with computational efficiency, as larger memory sizes show diminishing returns. SeTR’s performance proved equally intriguing, with a keep ratio of 20% (reducing representations by 80%) emerging as the optimal choice. Remarkably, even at extreme keep ratios as low as 5%, our model exhibits robust performance, with a very slight decrease in accuracy. Such results demonstrate the resilience of HERMES to hyperparameter variations suggesting that it can be deployed across diverse video understanding datasets with minimal hyperparameter tuning.
@@ 4.4 Qualitative Results
We present qualitative results on a challenging movie scene from the MovieChat-1k dataset to evaluate our model’s capability in answering both fine-grained and general questions about an extended video (14,000 frames). To rigorously assess the models, we bypass the original, basic Q&A from the dataset (e.g., Q: What’s the time in the video? A: Day; Q: Are there any shooting scenes? A: Yes) and pose questions that require a deeper understanding of the scene and actions, such as those concerning the expressions on the faces of the spectators. Our model demonstrates an ability to accurately respond to these questions while exhibiting a candid acknowledgment of its limitations (e.g., Q3). In contrast, MovieChat (Song et al., 2024) frequently generates hallucinated and incorrect answers. For instance, in response to Q2, MovieChat avoids answering the question directly and instead describes the scene. Notably, our model achieves this performance by processing only 100 out of the 14,000 frames (approximately 0.7%), whereas MovieChat processes 2,048 frames, which is more than 20 times the data utilized by our model. For additional qualitative results, including failure cases of our model, we refer the reader to section A.4.
@@ 6 Conclusion
We introduce HERMES: temporal-coHERent long-forM understanding with Episodes and Semantics, a novel framework designed to enhance long-form video understanding through two key components inspired by cognitive processes. The first, Episodic COmpressor (ECO), captures representations as sequences of continuous actions, reflecting episodic memory. The second, Semantics reTRiever (SeTR), serves as a high-level summarizer, distilling essential semantic information. Our model achieves state-of-the-art results on several long-video datasets, significantly outperforming existing methods. Through extensive experimentation, we have demonstrated the effectiveness and efficiency of ECO and SeTR.
@@ A.1 Reproducibility Statement
To facilitate the reproducibility of our work, we will make our code, pretrained models, default hyperparameters, and preprocessed annotations publicly available, with detailed hyperparameters for each dataset provided in Table 8. Our model demonstrates efficient performance, completing inference on the MovieChat-1k test set in 13 minutes (22 FPS) using a single V100 GPU (32 GB), and training on the MovieChat-1k dataset in less than 12 minutes with 8x 32 GB GPUs. In contrast to recent LLM-based approaches that necessitate extensive and costly multi-stage pretraining on increasingly large datasets, our model is designed for accessibility, thereby lowering the barrier for researchers without access to high-end computing resources. We achieve high performance while maintaining accessibility by leveraging existing pretrained weights and implementing our training-free ECO and SeTR, resulting in a model where finetuning is optional.
@@ A.2 Implementation Details
A.3 Extended Ablations
A.3.1 How does the number of frames affect the model’s accuracy and latency?
@@
A.3.2 How does the window size affect the model’s accuracy and latency?
@@
Latency exhibits a sharp decrease from window size 1 to 5, after which it remains relatively stable. This indicates that while smaller window sizes may seem computationally advantageous, they incur higher latency, possibly due to the need for more frequent ECO call and retrieval operations. The optimal trade-off occurs at a window size of 10, where we observe peak accuracy and stabilized latency suggesting that carefully tuned context windows can enhance long-form video understanding without incurring additional computational costs.
A.3.3 A Note on Latency
@@
A.4 More Qualitative Results
@@
A.5 Error Analysis: When does HERMES fail and why?
@@
A.6 How is our approach related to cognitive processes?

### v3

@@ 4.3 Ablation Studies
Table 7: Zero-shot performance comparison of MA-LMM with and without ECO and SeTR  integration on MovieChat-1k.
@@
We evaluate the impact of our Episodic COmpressor (ECO) as plug-ins to the existing MA-LMM model (He et al., 2024) by replacing the memory bank of MA-LMM. ECO is designed to efficiently process long video sequences while preserving temporal coherence and narrative structure. The results in Table 4.3 demonstrate substantial improvements with ECO integration with an accuracy increase by 3.4%. Moreover, ECO demonstrates superior efficiency compared to MA-LMM’s memory bank, almost halving the overall inference latency.
@@
To demonstrate the effectiveness and efficiency of our proposed ECO, we conduct a comparative analysis against two strong existing compression techniques: ToMe (Bolya et al., 2022) and MA-LMM (He et al., 2024) in Figure 5. We calculate the inference time for each model on the MovieChat-1k dataset. Powered by ECO our model achieves the highest accuracy (78.6%) among all models, outperforming MA-LMM by 5.3% and ToMe by a substantial 13.8%. HERMES achieves the highest inference speed among the compared models, while also maintaining superior accuracy. It is slightly faster than ToMe and significantly outperforms MA-LMM, reducing inference time by 46% compared to the latter. These results demonstrate our model’s ability to deliver state-of-the-art accuracy without compromising on efficiency.
@@
Our experiments on the MovieChat-1k (zero-shot) and Breakfast (fully-supervised) datasets reveal compelling insights into the optimal configuration of ECO (Figure 4) and SeTR (Figure 4). For ECO, we discover that an episodic memory size of 20 consistently yields peak performance across both datasets, achieving a 78.6% accuracy on MovieChat-1k and a 95.2% on Breakfast. This sweet spot balances comprehensive video representation with computational efficiency, as larger memory sizes show diminishing returns. SeTR’s performance proved equally intriguing, with a keep ratio of 20% (reducing representations by 80%) emerging as the optimal choice. Remarkably, even at extreme keep ratios as low as 5%, our model exhibits robust performance, with a very slight decrease in accuracy. Such results demonstrate the resilience of HERMES to hyperparameter variations suggesting that it is suitable for deployment across diverse video understanding datasets with minimal hyperparameter tuning.
@@ 4.4 Qualitative Results
We present qualitative results on a challenging movie scene from the MovieChat-1k dataset to evaluate our model’s capability in answering both fine-grained and general questions about an extended video (14,000 frames). To rigorously assess the models, we bypass the original, basic Q&A from the dataset (e.g., Q: What’s the time in the video? A: Day; Q: Are there any shooting scenes? A: Yes) and pose questions that require a deeper understanding of the scene and actions, such as those concerning the expressions on the faces of the spectators. Our model demonstrates an ability to accurately respond to these questions while exhibiting a candid acknowledgment of its limitations (e.g., Q3). In contrast, MovieChat (Song et al., 2024) frequently generates hallucinated and incorrect answers. For instance, in response to Q2, MovieChat avoids answering the question directly and instead describes the scene. Notably, our model achieves this performance by processing only 100 out of the 14,000 frames (approximately 0.7%), whereas MovieChat processes 2,048 frames, which is more than 20 times the data utilized by our model. For additional qualitative results, including failure cases of our model, please refer to Section A.5.
@@ 6 Conclusion
We propose HERMES, a novel framework designed to enhance long-form video understanding through two key components inspired by cognitive processes. The first, Episodic COmpressor (ECO), captures representations as sequences of continuous actions, reflecting episodic memory. The second, Semantics reTRiever (SeTR), serves as a high-level summarizer, distilling essential semantic information. Our model achieves state-of-the-art results on several long-video datasets, significantly outperforming existing methods. Through experiments on LVU, Breakfast, COIN and MovieChat, we have demonstrated the effectiveness and efficiency of ECO and SeTR.
@@ A.1 Reproducibility Statement
To facilitate the reproducibility of our work, we will make our code, pretrained models, default hyperparameters, and preprocessed annotations publicly available. Detailed hyperparameters for each dataset are also provided in Table 8. Our model demonstrates efficient performance, completing inference on the MovieChat-1k test set in 13 minutes (22 FPS) using a single V100 GPU (32 GB), and training on the MovieChat-1k dataset in less than 12 minutes with 8x 32 GB GPUs. In contrast to recent LLM-based approaches that necessitate extensive and costly multi-stage pretraining on increasingly large datasets, our model is designed for accessibility, thereby lowering the barrier for researchers without access to high-end computing resources. We achieve high performance while maintaining accessibility by leveraging existing pretrained weights and implementing our training-free ECO and SeTR, resulting in a model where finetuning is optional.
@@ A.2 Implementation Details
Figure 7: Illustration of our Episodic QFormer: We insert our ECO  in the original QFormer to effectively and efficiently compute and aggregate queries across long video sequences. It returns query episodes representing the whole video.
Figure 8: Illustration of SeTR: Our Semantics reTRiever  uses a stride of $k$ split the videos into groups $X$ of $N/k$ frames and $Y$ of $N-\frac{N}{k}$ frames, then merge each frame from $Y$ to its most semantically similar in $X$.
A.3 Model Details
A.3.1 Details of our Episodic QFormer
[…]
A.3.2 Details of SeTR
We design SeTR  as an efficient tool to retrieve semantic information from a long video. Given tokens extracted from a long video sequence, we use a stride of size $k$, to form a group of $\frac{N}{k}$ frames representing the number of semantics we want to extract. We then compress the remaining $N-\frac{N}{k}$ frames into extracted $\frac{N}{k}$ frames to obtain the semantic representations. SeTR  is illustrated in Figure 8.
A.4 Extended Ablations
A.4.1 How does the number of frames affect the model’s accuracy and latency?
@@
A.4.2 How does the window size affect the model’s accuracy and latency?
@@
Latency exhibits a sharp decrease from window size 1 to 5, after which it remains relatively stable. This indicates that while smaller window sizes may seem computationally advantageous, they incur higher latency, possibly due to the need for more frequent ECO call. The optimal trade-off occurs at a window size of 10, where we observe peak accuracy and stabilized latency suggesting that carefully tuned context windows can enhance long-form video understanding without incurring additional computational costs.
A.4.3 A Note on Latency
@@
A.5 More Qualitative Results
@@
A.6 Error Analysis: When does HERMES fail and why?
@@
A.7 How is our approach related to cognitive processes?